OpenStudy (anonymous):

Solve by undetermined coefficients/variation of parameters: \(x^3y'' + x^2y' -4xy = 1\) , x>0

OpenStudy (anonymous):

\[x^3y'' + x^2y' -4xy = 1\]\[x^2y'' + xy' -4y = \frac{1}{x}\]\[x^2y_c'' + xy_c' -4y_c =0\]\[\lambda (\lambda-1)+\lambda - 4 =0\]\[\lambda = \pm2\]\[y_c = c_1x^2+c_2x^{-2}\]

OpenStudy (anonymous):

hi @RolyPoly are you in DE

OpenStudy (anonymous):

\[u_1'(x^2)+u_2'(x^{-2})=0\]\[u_1'(2x)+u_2'(\frac{-2}{x^3})=\frac{1}{x}\] \[u_1'(x^2)+u_2'(x^{-2})=0\]\[u_1'(2x^2)+u_2'(\frac{-2}{x^2})=1\] \[4x^2u_1'=1\]\[u_1'= \frac{1}{4x^2}\]\[u_1 = -\frac{1}{4x}\] I'm doing something wrong, probably.. @mark_o. Yes.

OpenStudy (anonymous):

are you familiar with cauchy -euler equation? that is the problem you are having there

OpenStudy (anonymous):

are your class in that chapter already?

OpenStudy (anonymous):

chapter in cauchy -euler equation ?

OpenStudy (anonymous):

Yes. I'm having trouble with finding the particular solutions.

OpenStudy (anonymous):

The Cauchy-Euler part is done, I think.

OpenStudy (anonymous):

I divided both sides by x...

OpenStudy (anonymous):

And isn't it a second order DE only?

OpenStudy (anonymous):

Wwhat is it??

OpenStudy (anonymous):

thats the formula for getting the characteristic equation

OpenStudy (anonymous):

Why doesn't my method work?

OpenStudy (anonymous):

hold on lets double check it by using this formula

OpenStudy (anonymous):

And for Cauchy-Euler equation, isn't it the power of x and the no. of times of derivatives of y the same? (eg: x^2y'')

OpenStudy (anonymous):

it may not be the same so lets try doing the the formula one im giving you

OpenStudy (anonymous):

But this is not a 3rd order DE!

OpenStudy (anonymous):

im sorry i mess up,, yes its a 2nd order

OpenStudy (anonymous):

So will that m^3 thing work here?

OpenStudy (anonymous):

no,we have to use the (am^2+(b-a)m+c)x^m=0

OpenStudy (anonymous):

Wait.. Please refer to my first comment here. The Cauchy-Euler thing is done, I think.

OpenStudy (anonymous):

yes i think your ok

OpenStudy (anonymous):

The problem is how to find the particular solution.

OpenStudy (anonymous):

ok take x=e^t

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

so that we can reduced i/x into function of t

OpenStudy (anonymous):

Not i/x but 1/x

OpenStudy (anonymous):

yes 1/x

OpenStudy (anonymous):

and it's y=...(in terms of x) How come we have a t?!

OpenStudy (anonymous):

let x=e^t then 1/x==1/e^t=e^-t

OpenStudy (anonymous):

temporarily so that we can get Yp

OpenStudy (anonymous):

How come...

OpenStudy (anonymous):

assume Yp=Ae^-t

OpenStudy (anonymous):

It's not the form of yp

OpenStudy (anonymous):

Y=ae^-t Y'=-Ae^t Y''=Ae^t now sub this to your equation

OpenStudy (anonymous):

The yp is -1/(3x)...

OpenStudy (anonymous):

hmm try reading this first http://en.wikipedia.org/wiki/Cauchy%E2%80%93Euler_equation

OpenStudy (anonymous):

I think you need to know this is not a Cauchy-Euler equation.

OpenStudy (anonymous):

Since the right side \(\ne\) 0

OpenStudy (anonymous):

eventhough its not zero you can solve it using cauchy - euler eq

OpenStudy (anonymous):

this may be a good one to read http://www.sosmath.com/diffeq/second/euler/euler.html

OpenStudy (anonymous):

But Cauchy-Euler is over after we get the yc, isn't it?

OpenStudy (anonymous):

not until you convert everything to x=e^t

OpenStudy (anonymous):

try to convert the whole thing to x=e^t

OpenStudy (anonymous):

so let a=1, b=i,c=-4 then Y''(t)+0Y'(t)-4Y=e^-t that will be your equation

OpenStudy (anonymous):

I'm sorry, I must go now. Can we continue tomorrow?

OpenStudy (anonymous):

m^2_4=0 m=+2.-2 Yc(t)=C1e^-2t +C2e^2 next finding Yp

OpenStudy (anonymous):

that was m^2-4=0 m=-2,+2

OpenStudy (anonymous):

yes ok we can continue tomorrow... :D

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

ok YW good luck now and have fun :D

OpenStudy (anonymous):

@sirm3d can you give us some input here?

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

.......................... :D

OpenStudy (sirm3d):

by variation of parameters, let \(y=v\cdot\frac{1}{x^2}\)

OpenStudy (anonymous):

its been awhile since i took DE, i need to brush up again lol

OpenStudy (anonymous):

yesss variation of parameter too...but i think euler is much easier

OpenStudy (anonymous):

Looks like I'm not doing this right... \[y=\frac{v}{x^2}\]\[y'=-\frac{2v}{x^3}\]\[y''=-\frac{2v'}{x^3}+\frac{6v}{x^4}\] \[x^3(-\frac{2v'}{x^3}+\frac{6v}{x^4}) +x^2(-\frac{2v}{x^3}) - 4x(\frac{v}{x^2})=1\]\[-2v'+\frac{6v}{x} -\frac{2v}{x} - \frac{4v}{x}=1\]\[v'=-\frac{1}{2}\]\[v=-\frac{1}{2}x\] \[y_p=-\frac{1}{2x}\]Hmm... Where did I make mistakes?

OpenStudy (sirm3d):

differentiation of product yields \[\large y'=\frac{v'}{x^2}-2\frac{v}{x^3}\]

OpenStudy (anonymous):

But shouldn't we ''kill'' the last term to avoid troubles when differentiate y' again?

OpenStudy (anonymous):

Sorry, I meant the first term :\

OpenStudy (sirm3d):

\[\large y''=\frac{v''}{x^2}-2\frac{v'}{x^3}-2\frac{v'}{x^3}+6\frac{v}{x^4}\]

OpenStudy (anonymous):

Hmm.. Would you mind explaining why you did so? It's, somehow, different from what I've learnt in the lesson..

OpenStudy (sirm3d):

derivative of a product \(v\) and \(1/x^2\)

OpenStudy (anonymous):

I... mean why you did so...

OpenStudy (anonymous):

When way I've learnt is to let \(u_1y_1\), \(u_2y_2\)be the particular solution.. But somehow.. you're doing something different (in this question and in the previous one). I really want to know why you did that and how you came up with this y=v/x^2 ..

OpenStudy (sirm3d):

\[x^3y''+x^2y'-4xy=x^3\left( \frac{ v'' }{ x^2 }-4\frac{ v' }{ x^3 } +6\frac{ v }{ x^4 }\right)+x^2\left( \frac{ v' }{ x^2 }-\frac{ 2v }{ x^3 }\right)-4x \left( \frac{ v }{ x^2 } \right)\]\[=xv''-4v'+6\frac{ v }{ x }+v'-2\frac{ v }{ x }-4\frac{ v }{ x }\]\[=xv''-3v'=1\]

OpenStudy (sirm3d):

the idea behind the variation of parameter is to use one complementary function that is a solution of the homogeneous DE. In this problem, you produced one complementary function\(1/x^2\) using cauchy-euler method.

OpenStudy (sirm3d):

the variation of parameter introduces a function \(v\) to solve THE particular solution using one complementary solution \(1/x^2\) by assuming that \(y_p=v(1/x^2)\) is the particular solution of the DE, where v is to be determined.

OpenStudy (sirm3d):

\[\large xv''-3v'=1\]let \(w=v'\) and \(w' = v''\)\[\large xw'-3w=1\]\[\large w'-\frac{3}{x}w=\frac{1}{x}\]

OpenStudy (sirm3d):

\[\large w(\frac{ 1 }{ x^3 })=\int\limits \frac{1}{x^4} dx=-(1/3)x^{-3}\]\[\large v'=w=(-1/3)\]\[\large v=(-1/3)x\]\[\large y_p=v(\frac{ 1 }{ x^2 })=-\frac{ 1 }{ 3x }\]

OpenStudy (sirm3d):

check: \[\large x^3y''+x^2y'-4xy=x^3\left( -\frac{ 2 }{ 3x^3 } \right)+x^2\left( \frac{ 1 }{ 3x^2 } \right)-4x \left( -\frac{ 1 }{ 3x } \right)\]\[\large -\frac{ 2 }{ 3 }+\frac{ 1 }{ 3 }+\frac{ 4 }{ 3 }=1\] and so it's solved.

OpenStudy (anonymous):

But I think there are two complementary solutions? \(x^2\) and \(x^{-2}\)?!

OpenStudy (sirm3d):

variation of parameter allows you to choose any of the complementary solutions. i chose \(x^{-2}\) over \(x^2\) for no particular reason. why don't you try \(y=vx^2\) ? i have yet to learn your method to find why you get -1/4 when i got -1/3

OpenStudy (anonymous):

Wouldn't it be extremely weird?! I think I chose both the complementary functions..but you just chose one.. It's weird...

OpenStudy (anonymous):

Consider y''-4y'+4y = e^(2x) / x \[y_c = c_1e^{2x} +c_2xe^{2x}\] The way I find the particular solution is \[u_1'e^{2x}+u_2'(xe^{2x})=0\]\[u_1'(2e^{2x})+u_2'(2xe^{2x}+e^{2x})=\frac{e^{2x}}{x}\]and find u1 u2. Then, there are two parts in the particular solution...

OpenStudy (sirm3d):

is the particular solution \[\large y_p=-x e^{2x}+(\ln x) x e^{2x}\]

OpenStudy (sirm3d):

i got that result following your method.

OpenStudy (sirm3d):

if i use my method, and choose the complementary solution \(e^{2x}\)\[\large y=v e^{2x}\]\[\large y'=v'e^{2x}+2v e^{2x}\]\[\large y''=v'' e^{2x}+4v'e^{2x}+4v e^{2x}\]

OpenStudy (sirm3d):

\[\large y''-4y' + 4y=(v''e^{2x}+4v'e^{2x}+4v e^{2x})-4(v'e^{2x} + 2v e^{2x})+4(v e^{2x})\]\[\large =v''e^{2x}=\frac{e^{2x}}{x}\]\[\large v''=\frac{1}{x}\]

OpenStudy (sirm3d):

\[\large v'=\ln x\]\[\large v=x\ln x - x\]\[\large y=\left(x \ln x - x \right) e^{2x}\], no different from the result using your method.

OpenStudy (anonymous):

Is.. it...magic? :S Please let me try to do it using your method!!

OpenStudy (sirm3d):

can you provide me a link on your method?

OpenStudy (anonymous):

Link!? I don't have it :( Or can I just type the notes I copied in the lesson here? (scanner is not working now :( )

OpenStudy (sirm3d):

i found a link detailing your method. i'm reading it now.

OpenStudy (sirm3d):

@RolyPoly i saw some mistakes in your solution. your two equations are \[\large u_1'x^2+u_2x^{-2}=0\]\[\large u_1(2x)+u_2(-2x^{-3})=\frac{1/x}{x^2}=\frac{1}{x^3}\]

OpenStudy (anonymous):

Why should it be (1/x)/x^2 ??

OpenStudy (sirm3d):

just solve \(u_1\) and \(u_2\) and sub them in \[\large y_p= u_1x^2+u_2x^{-2}\] and you should get the same answer \(-1/(3x)\)

OpenStudy (sirm3d):

the second equation in the system is of the form \[\large u_1'y_1'+u_2'y_2'=\frac{r(x)}{b_nx^n}\] from the cauchy-euler \[\large b_nx^n y^{(n)}+\cdots +b_1xy'+b_0y=r(x)\]

OpenStudy (sirm3d):

you should get \(y_p=-1/(3x)\) this time after solving \(u_1\) and \(u_2\). this are the values i got: \[\large u_1=-\frac{1}{12x^3}\]\[\large u_2=-\frac{x}{4}\]

OpenStudy (sirm3d):

@RolyPoly hope i have given you peace of mind in this problem.

OpenStudy (anonymous):

My... teacher... didn't .... told... us... that... the ...second equation is in the form \[ u_1'y_1'+u_2'y_2'=\frac{r(x)}{b_nx^n}\]

OpenStudy (sirm3d):

that may be the cause of your problems.

OpenStudy (anonymous):

Most probably.. That seems your method works better.. I have to try both methods to decide which to use.. Thanks!!! I hope you don't mind me tagging you if I still have problems concerning this question!